OpenStudy (solomonzelman):

I must be tired or stupid, but this problem did confuse me. Anyone kind enough to help?

OpenStudy (solomonzelman):

A population in place A is 50,000 in January 1, 2010. Every year the population in place A increases by 7%.. The population in place B is 70,000 in January 1, 2010. Every year the population in place A decreases by 5%. at what year will the population of place A exceed the population in place B? (slow typer, sorry)

OpenStudy (solomonzelman):

smart A S S ... I never thought of googling math questions.

OpenStudy (the_fizicx99):

Wouldn't you use the geometric formula? f(x)=P(1+r)^x

OpenStudy (blackops2luvr):

BA BOOM

OpenStudy (anonymous):

Is the answer 18?

OpenStudy (solomonzelman):

I admit I suck at math, but not that much.

OpenStudy (anonymous):

Solve \[ 50000 (1+0.07)^t=70000 (1+.05)^t \] for t

OpenStudy (the_fizicx99):

O_O noooo, not you

OpenStudy (solomonzelman):

I already did, and I got 3 as my answer for t.

OpenStudy (solomonzelman):

oops

OpenStudy (solomonzelman):

Actually after 4 years place A will be greater.

OpenStudy (the_fizicx99):

Did you miss out that it says A decreases? 1.05 is a growth rate (above 0.99) or I'm i just confused ?

OpenStudy (the_fizicx99):

am I *

OpenStudy (solomonzelman):

idk abt u, but ik I found the answer to my problem.

OpenStudy (the_fizicx99):

._. something like this? >.> idk either

OpenStudy (solomonzelman):

place B is decreasing by 7%. not increasing.

OpenStudy (solomonzelman):

Not by 7, by 5, but it is DEcreasing.

OpenStudy (the_fizicx99):

I see where i might've gone wrong >.>

OpenStudy (solomonzelman):

okay, that was a good brain exercise, and thx everyone!

OpenStudy (the_fizicx99):

-sigh- bo